calculate ecell for the following electrochemical cell at 25 C.

Pt(s)| H2 (g,1.00 atm) | H+ (aq, 1.00 M) || (Pb2+ (aq, 0.150 M) | Pb (s)

Pb^2+(aq) + 2e- ---> Pb (s) Eo=-0.126 V

2H^+ (aq) + 2e- ---> H2 (g) Eo=0.00 V

Thank you!
Answer #1

eesoade = 0.00-(-0.126)V = 0.126V 0.0592 v, [Pb2 log 0.0592 V 0.150 = 0.126-ww-log1002 = 0.102 V